# MSDiagnosis: An EMR-based Dataset for Clinical Multi-Step Diagnosis

### v1

@@ Abstract
Clinical diagnosis is critical in medical practice, typically requiring a continuous and evolving process that includes primary diagnosis, differential diagnosis, and final diagnosis. However, most existing clinical diagnostic tasks are single-step processes, which does not align with the complex multi-step diagnostic procedures found in real-world clinical settings. In this paper, we propose a multi-step diagnostic task and annotate a clinical diagnostic dataset (MSDiagnosis). This dataset includes primary diagnosis, differential diagnosis, and final diagnosis questions. Additionally, we propose a novel and effective framework. This framework combines forward inference, backward inference, reflection, and refinement, enabling the LLM to self-evaluate and adjust its diagnostic results. To assess the effectiveness of our proposed method, we design and conduct extensive experiments. The experimental results demonstrate the effectiveness of the proposed method. We also provide a comprehensive experimental analysis and suggest future research directions for this task.
@@
MSDiagnosis: An EMR-based Dataset for Clinical Multi-Step Diagnosis
@@ 1 Introduction
Figure 1: An example of our diagnostic dataset and its differences from the previous diagnostic dataset. The doctor first makes a primary diagnosis based on the chief complaint, examination, and medical history, then determines the differential disease, and finally makes a final diagnosis based on the hospital course.
Electronic Medical Records (EMRs) digitally document comprehensive information about the entire diagnostic and treatment process of a patient Bates et al. (2003). This encompasses clinical data such as hospital courses, clinical notes, and diagnostic results Keyhani et al. (2008). The clinical notes in the EMRs are medical texts written by doctors during consultations to address the patient’s medical history, chief complaints, and examinations. Through these records, doctors can use their expertise to make a diagnosis for the patient Tomašev et al. (2021).
Dataset Data Source Q.Type w/o Criteria w/o Multi-step. Size
@@
Table 1: Overview of the clinical diagnosis dataset. “*” indicates that the dataset contains data for multiple tasks, with the diagnosis being just one of them. In the “Size” column, the value outside the brackets indicates the number of samples related to diagnosis, and the value inside the brackets indicates the total number of samples in the dataset. The terms “Criteria”, “Multi-step”, “Q.Type”, and “MCQ” stand for “diagnostic criteria”, “multi-stage diagnosis”, “question type”, and “multiple-choice QA” respectively.
@@
The clinical diagnosis process is a continuous and evolving process Tiffen et al. (2014). Specifically, doctors make primary and differential diagnoses based on the patient’s medical history, chief complaint, and laboratory and aided examination. The final diagnosis is then determined by considering clinical changes throughout the diagnosis and treatment process. However, a significant gap remains between the existing diagnostic dataset and actual clinical practice. Most existing diagnoses are single-step processes Wang et al. (2024) Lyu et al. (2023) Liu et al. (2024a), where a diagnosis is made directly based on the patient’s medical history, chief complaint, and examination results. This single-step approach does not align with the multi-step process typically used in clinical work.
@@
Hence, in this paper, we propose a multi-step diagnostic task and annotate a clinical diagnosis dataset called ReaDiagnosis. In the multi-step diagnostic task, a primary diagnosis and diagnostic criteria are generated based on the patient’s medical history, chief complaint, and other relevant information. Then, a differential diagnosis is made to narrow down the possible diseases. Finally, the final diagnosis and diagnostic criteria are made by combining the hospital course, primary diagnosis, and differential diagnosis. The specific process is shown in Fig. 1. Additionally, our ReaDiagnosis dataset is designed to evaluate the diagnostic capabilities of LLMs. The dataset comprises 2,225 medical records collected from medical websites, covering 12 departments. Each medical record contains five related diagnostic questions. For ease of comparison, we list the most relevant works in Table 1. To accurately assess diagnostic criteria, we annotate each answer with corresponding scoring points and evaluate the quality of the answers by calculating the macro-recall for each scoring point.
@@
For the multi-step diagnostic task, we propose a simple and effective pipeline framework. This framework consists of two stages. The first stage is forward inference. Specifically, we first combine LLM for forward inference diagnosis by emphasizing key information in EMRs and merging similar EMRs. The second stage involves backward inference and reflective optimization. Specifically, we design a reflective optimization strategy and a backward inference strategy that allow the LLM to refine its diagnostic results.
In summary, the contributions of our study are as follows:
  • We propose a multi-step diagnostic task and annotate a high-quality dataset (ReaDiagnosis). This dataset contains 2,225 cases, each designed with questions corresponding to multi-stage diagnosis.
  • We introduce a novel and efficient framework that integrates forward and backward inference, enabling LLMs to perform bidirectional reasoning for validation, thereby achieving reflective refinement.
@@
  • We conduct extensive experiments, including using open-source and closed-source LLMs, manual evaluation, and the framework proposed in this paper, to comprehensively assess performance on multi-step diagnostic tasks. The experimental results demonstrate the effectiveness of the proposed method.
@@ 2 Problem Formulation
In this paper, we consider the multi-step diagnosis task as a multi-round dialogue problem. For each complex EMR, we simulate the interaction between an examiner and a candidate, where the candidate is required to provide an answer $A$ to a specific diagnostic question $Q$. Specifically, given a patient’s admission record $E$, which includes the chief complaint, present history, past history, physical examination, and laboratory and aided examination, the candidate answers two questions $R_{1}=[Q_{1},Q_{2}]$ in the first round of dialogue. These questions mainly involve the patient’s primary diagnosis and its diagnostic criteria. As illustrated in Fig. 3, $Q_{1}$ and $Q_{2}$ are “What is the patient’s primary diagnosis?” and “What is the criterion for the primary diagnosis?” respectively. In the second round of dialogue, the candidate needs to answer the question $R_{2}=[Q_{3}]$. This question refers to asking the patient about their differential diagnosis. In the third round of dialogue, the candidate receives the patient’s hospital course $T$, the dialogue history $H$, and two additional questions $R_{3}=[Q_{4},Q_{5}]$. These questions focus on the patient’s final diagnosis and its diagnostic criteria. Therefore, the input $X$ for the current dialogue can be formally expressed as $X=E+T+H+Q_{i}$, where $Q_{i}\in R_{3}$. The reference answer for this input is $A_{i}$. Note that the question may be phrased differently (see Section 3.2.1 for explanations).
3 Data Construction
In this section, we detail the construction process of the ReaDiagnosis dataset, including the data collection, data pre-processing, and data annotation.
3.1 Data Collection and Pre-processing
In this study, we select a Chinese medical website [^1] https://www.iiyi.com/ as the source of EMRs. This website is an open-source medical platform that contains a large number of real EMRs. After de-identifying the data, we obtain a total of 11,900 EMRs. To ensure high-quality EMR, we preprocess the data through three steps. First, we remove redundant information unrelated to medical content, such as HTML web elements, copyright statements, and advertisements. Second, we eliminate EMRs with duplicate field values and missing fields, such as those lacking primary diagnosis, differential diagnosis, final diagnosis, or hospital course fields. Third, we deduplicate EMRs from the same department. If two records have identical chief complaints, present histories, and physical examinations, one of them is removed. After data preprocessing, we ultimately obtain 3,501 high-quality, complex, and authentic EMRs.
@@ 3.2.1 Question Construction
We first manually construct five questions, including the patient’s primary diagnosis, the criterion for the primary diagnosis, the differential diagnosis, the final diagnosis, and the criterion for the final diagnosis. The definition of each question is shown in Table 7 in Appendix D. Then, we use GPT-4 to expand each constructed question with ten similar questions. Finally, we manually check and filter out unreasonable question expressions. When constructing EMR questions, we randomly select from these similar questions to enhance the diversity of the questions.
@@ 3.2.2 Answer Annotation
To ensure the quality of the dataset construction, we form a professional team comprising three inspectors and one reviewer. All team members undergo specialized medical training to understand primary diagnosis, differential diagnosis, and final diagnosis. The construction process includes three stages: the first round of annotation, the second round of inspection, and the third round of review.
The first round of annotation. For the primary diagnosis, differential diagnosis, and final diagnosis, we directly take the values from the corresponding fields in the EMR as the standard answers for the respective diagnostic questions. For the criterion for the primary and final diagnosis, we use GPT-4 to generate these diagnostic criteria.
The second round of inspection. We engage three university students to simultaneously check the rationality of all question-answer pairs. Samples that all three inspectors considered unreasonable are directly discarded. If one or two inspectors find a sample unreasonable, it is manually re-annotated and retained only if all three inspectors subsequently agree on its reasonableness.
The third round of review. A verified batch is given to a senior doctor for double review. Medical experts randomly check 20% of the batch. Unqualified annotations are returned to the inspection team, which can further refine the standards, thus standardizing the construction team’s work. This process is repeated until the batch accuracy reaches 95%.
3.2.3 Scoring Point Annotation
To more accurately evaluate open-ended questions such as primary diagnostic criteria and final diagnostic criteria, we annotate the score point of the answers to these questions. Specifically, for each primary and final diagnostic criteria, we first construct score point extraction prompts and employ GPT-4 to extract the information from the standard answers. The score point for the diagnostic criteria mainly includes four categories: medical history, symptoms, physical signs, and test results. Then, we invite two students with specialized medical training to validate the extracted information. If any inconsistencies are found in the EMRs, the data is re-annotated. If a sample remains unreasonable after three attempts, it is discarded. Finally, we select the top 2,225 high-quality samples to form the ReaDiagnosis. We calculate the Cohen’s Kappa Banerjee et al. (1999) score for the key information of the primary diagnostic criteria and the final diagnostic criteria. The Cohen’s Kappa for the primary diagnostic criteria is 0.81, indicating a high level of agreement, while the Cohen’s Kappa for the final diagnostic criteria is 0.79, indicating a moderate level of agreement.
@@ 4 Dataset Analysis
In this section, we introduce the statistics of ReaDiagnosis and the characteristics of the ReaDiagnosis in detail.
@@ 4.1 Data statistics
Figure 2: Distribution of “Primary diagnosis matches final diagnosis” and “Primary diagnosis differs from final diagnosis” in ReaDiagnosis across different departments.
Type Size Diag-C Diag-D Avg1 Avg2
@@
Table 2: The statistics of our constructed ReaDiagnosis dataset. “Diag-C” refers to EMRs where the primary diagnosis matches the final diagnosis, and “Diag-D” refers to EMRs where the primary diagnosis differs from the final diagnosis. “Avg1” indicates the average number of diseases identified in the primary diagnosis, and “Avg2” indicates the average number of diseases identified in the final diagnosis.
As reported in Table 2, the ReaDiagnosis dataset contains a total of 2,225 EMRs covering 12 departments, which are divided into training, validation, and test sets according to a 7:1:2 ratio. Additionally, we conduct a detailed statistical analysis of patients’ diagnoses from three main perspectives. 1) EMRs where the primary diagnosis matches the final diagnosis, and EMRs where the primary diagnosis differs from the final diagnosis. Fig. 2 shows a detailed visualization of the statistical data. In the category where the primary and final diagnoses are the same, surgery has the highest proportion (approximately 48.02%). In the category where the primary and final diagnoses are different, internal medicine has the highest proportion (approximately 31.64%). 2) Number of Diseases: The average number of diseases diagnosed at the primary diagnosis stage is 2, with a maximum of 21 diseases. At the final diagnosis stage, the average number of diseases is 3, with a maximum of 21 diseases. 3) Types of diagnostic changes: Through statistical analysis of diagnosis changes, we find that changes mainly fall into three categories: addition, deletion, and modification.
Figure 3: Our framework for the multi-step clinical diagnosis. The top portion of the figure illustrates the flow of the framework, comprising two stages. The first stage involves the forward inference diagnosis. The second stage focuses on backward inference and reflection.
@@ 4.2 Data Characteristics
The ReaDiagnosis dataset has several significant advantages compared to previous medical diagnostic datasets: 1) Reliability and Authenticity of Data: The ReaDiagnosis dataset is entirely based on real patient EMRs, including detailed treatment plans and course records. This real clinical data source ensures high reliability and authenticity, which helps simulate the actual medical environment more accurately. 2) Rich Variety of Departments and Diseases: The ReaDiagnosis dataset covers patient EMRs from 12 different departments and includes various types of diseases, such as common diseases, rare diseases, acute diseases, and chronic diseases. This wide coverage of departments and diverse disease types ensures the broad applicability of the data. 3) Inclusion of Multi-step Diagnostic Processes: This multi-step diagnostic process is more in line with actual clinical diagnostic scenarios. By recording diagnostic changes, the dataset better reflects the complex diagnostic paths and dynamic changes in real-world medical practice.
5 Methods
@@ 5.1 Framework
As illustrated in Fig. 3, our proposed framework mainly consists of two stages. The first stage involves forward inference. In this stage, we retrieve similar EMRs to serve as in-context learning (ICL) for the diagnosis, allowing the LLM to diagnose the patient. The second stage is backward inference and reflection. In this stage, we first validate the diagnostic criteria against the facts derived from the diagnostic results. Then, the LLM uses designed reflection rules to reflect on the diagnostic outcomes. Finally, the diagnosis is optimized by integrating all the previous results.

### v3

# MSDiagnosis: A Benchmark for Evaluating Large Language Models in Multi-Step Clinical Diagnosis

@@ Abstract
Clinical diagnosis is critical in medical practice, typically requiring a continuous and evolving process that includes primary diagnosis, differential diagnosis, and final diagnosis. However, most existing clinical diagnostic tasks are single-step processes, which does not align with the complex multi-step diagnostic procedures found in real-world clinical settings. In this paper, we propose a Chinese clinical diagnostic benchmark, called MSDiagnosis. This benchmark consists of 2,225 cases from 12 departments, covering tasks such as primary diagnosis, differential diagnosis, and final diagnosis. Additionally, we propose a novel and effective framework. This framework combines forward inference, backward inference, reflection, and refinement, enabling the large language model to self-evaluate and adjust its diagnostic results. To this end, we test open-source models, closed-source models, and our proposed framework. The experimental results demonstrate the effectiveness of the proposed method. We also provide a comprehensive experimental analysis and suggest future research directions for this task.
@@
MSDiagnosis: A Benchmark for Evaluating Large Language Models in Multi-Step Clinical Diagnosis
@@ 1 Introduction
Figure 1: An example of our diagnostic benchmark and its differences from the previous diagnostic benchmark.
Clinical diagnosis is a central element of clinical decision-making, involving the integration of chief complaint, present history, and physical examinations, along with clinical experience and medical knowledge, to make a scientifically grounded judgment regarding the nature and cause of a disease Ball et al. (2015). Accurate diagnosis helps to provide the most appropriate treatment for patients. In recent years, research in data-driven clinical diagnosis has advanced rapidly, particularly with the widespread use of electronic medical records (EMRs), which enable diagnostic models to integrate multi-stage patient data for more precise analysis Herrero-Zazo et al. (2021).
Table 1: Overview of the clinical diagnosis benchmark. “*” indicates that the dataset contains data for multiple tasks, with the diagnosis being just one of them. In the “Size” column, the value outside the brackets indicates the number of samples related to diagnosis, and the value inside the brackets indicates the total number of samples in the dataset. The terms “Criteria”, “Multi-Step”, “Q.Type”, and “MCQA” stand for “diagnostic criteria”, “multi-step diagnosis”, “question type”, and “multiple-choice QA” respectively.
Dataset Data Source Q.Type Criteria Multi-Step Key Points Size
@@
The clinical diagnosis process is a continuous and evolving process Tiffen et al. (2014). Specifically, doctors first make a primary diagnosis based on the patient’s chief complaint, present history, and laboratory and aided examination. Then they narrow down the diagnostic options through differential diagnosis. Finally, they determine the final diagnosis by considering clinical changes throughout the diagnostic and treatment process. However, a significant gap remains between the existing diagnostic benchmark and actual clinical practice. Most existing diagnostic tasks are single-step processes Wang et al. (2024); Lyu et al. (2023); Liu et al. (2024a), where a diagnosis is made directly based on the patient’s medical history, chief complaint, and examination results. This single-step approach does not align with the multi-step process typically used in clinical practice. In addition, current evaluations of single-step diagnostic tasks typically use BLEU and ROUGE metrics. However, these metrics, which rely solely on lexical overlap, cannot effectively assess the rationale behind the diagnosis or the coverage of key evidence.
@@
Hence, in this paper, we propose a Multi-Step clinical Diagnosis benchmark (MSDiagnosis). In this benchmark, a primary diagnosis and diagnostic criteria are generated based on the patient’s chief complaint, present history, and examination results. Then, a differential diagnosis is made to narrow down the possible diseases. Finally, the final diagnosis and diagnostic criteria are made by combining the hospital course, primary diagnosis, and differential diagnosis. The specific process is shown in Fig. 1. The MSDiagnosis benchmark is designed to evaluate the diagnostic capabilities of large language models (LLMs). The benchmark comprises 2,225 medical records collected from medical websites, covering 12 departments. Each medical record contains five related diagnostic questions. For ease of comparison, we list the most relevant works in Table 1. Additionally, to accurately assess diagnostic criteria, we annotate each answer with corresponding key points and evaluate the quality of the answers by calculating the macro-recall for each key point.
@@
For the multi-step clinical diagnostic benchmark, we propose a simple and effective pipeline framework. This framework consists of two stages. The first stage is forward inference. Specifically, we retrieve similar EMRs to serve as in-context learning (ICL), allowing the LLM to diagnose the patient. The second stage involves backward inference, reflection, and refinement. Specifically, we design a backward inference strategy, a reflection strategy, and a refinement strategy, enabling the LLM to check and refine its diagnostic results.
Our contributions are summarized as follows:
  • We propose a multi-step clinical diagnosis benchmark (MSDiagnosis) that includes three tasks: primary diagnosis, differential diagnosis, and final diagnosis.
  • We propose a simple and efficient framework that combines forward and backward inference, enabling LLMs to validate and refine diagnostic results.
@@
  • We evaluate both open-source and closed-source LLMs on MSDiagnosis and conduct extensive experiments on our framework. The experimental results demonstrate the effectiveness of the proposed method.
@@ 2 Problem Formulation
In this paper, we consider the multi-step diagnosis task as a multi-round dialogue problem. For each complex EMR, we simulate the interaction between an examiner and a candidate, where the candidate is required to provide an answer to a specific diagnostic question. Specifically, given a patient’s admission record $E$, which includes the chief complaint, present history, past history, physical examination, and laboratory and aided examination, the candidate answers two questions $R_{1}=[Q_{1},Q_{2}]$ in the first round of dialogue. These questions mainly involve the patient’s primary diagnosis and its diagnostic criteria. As illustrated in Fig. 3, $Q_{1}$ and $Q_{2}$ are “What is the patient’s primary diagnosis?” and “What is the criterion for the primary diagnosis?” respectively. In the second round of dialogue, the candidate needs to answer the question $R_{2}=[Q_{3}]$. This question refers to asking the patient about their differential diagnosis. In the third round of dialogue, the candidate receives the patient’s hospital course $T$, the dialogue history $H$, and two additional questions $R_{3}=[Q_{4},Q_{5}]$. These questions focus on the patient’s final diagnosis and its diagnostic criteria. Therefore, the input $X$ for the current dialogue can be formally expressed as $X=E+T+H+Q_{i}$, where $Q_{i}\in R_{3}$. The reference answer for this input is $A_{i}$.
3 MSDiagnosis
In this section, we detail the construction process of the dataset, including the data collection, selection, and data annotation.
3.1 Data Collection and Selection
In this study, we select a Chinese medical website [^1] https://www.iiyi.com/ as the source of EMRs. The raw data are anonymized, and we apply additional anonymization to ensure no protected health information (PHI) remains. After de-identifying the data, we obtain a total of 11,900 EMRs. To ensure high-quality EMR, we select the data through four steps. First, we remove redundant information unrelated to medical content, such as web elements, copyright statements, and advertisements. Second, we eliminate EMRs with duplicate field values and missing fields, such as those lacking primary diagnosis, differential diagnosis, final diagnosis, or hospital course fields. After this step, we retain 5096 EMRs. Third, we deduplicate EMRs from the same department. If two records have identical chief complaints, present histories, and physical examinations, one of them is removed. After this step, we retain 4179 EMRs. Fourth, to avoid duplicate patients in the dataset, we conduct a meticulous matching process using patients’ demographic information (such as gender, occupation, etc.) and filter out individuals with identical profiles. After rigorous screening, we ensure the dataset contains no patients with fully identical records. After four steps, we ultimately obtain 3,501 high-quality, complex, and authentic EMRs.
@@ 3.2.1 Question Construction
We first manually construct five seed questions, including the patient’s primary diagnosis and its criteria, the differential diagnosis, and the final diagnosis and its criteria. The definition of each question is shown in Appendix D. Then, we use GPT-4 to expand each constructed question with ten similar questions. Finally, we manually check and filter out unreasonable question expressions.
@@ 3.2.2 Answer Annotation
To ensure the quality of the dataset construction, we form a professional team comprising three inspectors and one reviewer. All team members undergo specialized medical training to understand primary diagnosis, differential diagnosis, and final diagnosis. The construction process includes three stages: first-round annotation, second-round checking, and third-round review.
First-round annotation. For questions related to initial diagnosis, differential diagnosis, and final diagnosis, if the original medical records contain relevant answers, the standard answers are directly extracted from the EMR. However, statistical analysis reveals that 1,989 cases lack diagnostic results. Additionally, since the original data does not provide criteria for the primary and final diagnoses, we initially use GPT-4 to generate diagnostic criteria for 3,501 cases.
Second-round checking. We engage three university students to simultaneously check the rationality of all question-answer pairs. Samples that all three inspectors considered unreasonable are directly discarded. If one or two inspectors find a sample unreasonable, it is manually re-annotated and retained only if all three inspectors subsequently agree on its reasonableness. After rigorous check, we identify 1,222 cases that are unanimously deemed unreasonable.
Third-round review. A verified batch is given to a medical expert for double review. The medical expert randomly inspects 20% of the batch samples. Any unqualified annotations are returned to the check team with explanations, which can further refine the standards. This process is repeated until the batch accuracy reaches 95%. After this stage, we remove 54 cases, ultimately retaining 2,225 high-quality cases.
3.2.3 Key Points Annotation
To more accurately evaluate open-ended questions such as primary diagnostic criteria and final diagnostic criteria, we annotate the key points of the answers to these questions. Specifically, for each primary and final diagnostic criteria, we first construct key point extraction prompts and employ GPT-4 to extract the information from the standard answers. The key point for the diagnostic criteria mainly includes four categories: medical history, symptoms, physical signs, and examination results. Then, we invite two students with specialized medical training to validate the extracted information. If any inconsistencies are found in the EMRs, the data is re-annotated. Finally, we obtain 2,225 high-quality samples to construct MSDiagnosis. We calculate the Cohen’s Kappa Banerjee et al. (1999) score for the key points of the primary diagnostic criteria and the final diagnostic criteria. The Cohen’s Kappa for the primary diagnostic criteria is 0.81, indicating a high level of agreement, while the Cohen’s Kappa for the final diagnostic criteria is 0.79, indicating a moderate level of agreement.
@@ 4 Dataset Analysis
In this section, we introduce the statistics and the characteristics of the MSDiagnosis in detail.
@@ 4.1 Data Statistics
Figure 2: Distribution of “Primary diagnosis matches final diagnosis” and “Primary diagnosis differs from final diagnosis” in MSDiagnosis across different departments.
Table 2: The statistics of our constructed dataset. “Diag-M” and “Diag-D” refer to EMRs where the primary diagnosis is the same as or different from the final diagnosis, respectively. “AvgP” and “AvgF” represent the average number of diseases in the primary and final diagnoses, respectively.
Type Size Diag-M Diag-D AvgP AvgF
@@
As reported in Table 2, the dataset contains a total of 2,225 EMRs covering 12 departments, which are divided into training, validation, and test sets according to a 7:1:2 ratio. Additionally, we conduct a detailed statistical analysis of patients’ diagnoses from three main perspectives. 1) Department Distribution. Fig. 2 presents the distribution of EMRs across different departments, categorized by whether the primary diagnosis is consistent with the final diagnosis. In the category where the primary and final diagnoses are the same, surgery has the highest proportion (48.02%). In the category where the primary and final diagnoses are different, internal medicine has the highest proportion (31.64%). 2) Number of Diseases. The average number of diseases diagnosed in the primary diagnosis stage is 2, with a maximum of 21 diseases. At the final diagnosis stage, the average number of diseases is 3, with a maximum of 21 diseases. 3) Types of Diagnostic Changes. Statistical analysis reveals that diagnosis changes primarily fall into three categories: addition, deletion, and modification.
Figure 3: Our framework for the multi-step clinical diagnosis. The top portion of the figure illustrates the flow of the framework, comprising two stages. The first stage involves the forward inference diagnosis. The second stage focuses on backward inference, reflection, and refinement.
@@ 4.2 Data Characteristics
The MSDiagnosis has several significant advantages compared to previous medical diagnostic datasets: 1) Reliability and Authenticity of Data. The MSDiagnosis is entirely based on EMRs, including detailed treatment plans and course records. 2) Rich Variety of Departments and Diseases. The MSDiagnosis covers patient EMRs from 12 different departments and includes various types of diseases, such as common diseases, rare diseases, acute diseases, and chronic diseases. This wide coverage of departments and diverse disease types ensures the broad applicability of the data. 3) Inclusion of Multi-step Diagnostic Processes. This multi-step diagnostic process is more in line with actual clinical diagnostic scenarios. By recording diagnostic changes, the dataset better reflects the complex diagnostic paths and dynamic changes in real-world medical practice.
5 Method
@@ 5.1 Framework
As illustrated in Fig. 3, our proposed framework mainly consists of two stages. The first stage involves forward inference. In this stage, we retrieve similar EMRs to serve as ICL, enabling the LLM to diagnose the patient. The second stage is backward inference, reflection, and refinement. In this stage, we first validate the diagnostic criteria against the facts derived from the diagnostic results. Then, the LLM uses designed reflection rules to evaluate the diagnostic outcomes. Finally, the diagnosis is refined by integrating all the previous results.